### v2

# A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor

Abstract
The strict geometric rules that define aperiodic tilings lead to the unique spectral and transport properties of quasicrystals, but also limit our ability to design them. In this work, we explore the first continuously tunable family of two-dimensional aperiodic tilings in which the underlying real-space geometry becomes a control knob of the wave-function’s quantum geometric tensor. The real-space geometry can be used to tune into topological phases occupying an expanded phase space compared to crystals, or into a disorder-driven topological Anderson insulator. The quantum metric can also be tuned continuously, opening new routes towards tunable single- and many-body physics in aperiodic solid-state and synthetic systems.
Introduction—The electronic properties of quasicrystals can be strikingly different to those of crystals [1, 2, 3]. For example, quasicrystals exhibit rotational symmetries forbidden in crystals, and their wave-functions are neither localized nor extended [1, 2, 3]. However, the strict mathematical rules that determine the quasicrystalline atomic arrangements and the special wave-functions they host also restricts the possibility of tuning them. Different realizations of two-dimensional quasicrystals, such as the Penrose [4] or the Ammann-Beenker [5] tilings, cannot be smoothly deformed into one another. Hence, continuously tuning the electronic properties of quasicrystals is typically limited to changing model parameters, as in crystals [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
Figure 1: Tiles and tilings for different values of the tile side length, defined by two side lengths, $\ell$ and $1-\ell$. (a) shows a selection of tiles, including the Chevron, Hat, Turtle, and Comet. Their corresponding aperiodic tilings are shown in (b-e). For each $\ell$ we place two orbitals $l$ at every vertex, with onsite terms $\epsilon_{ll^{\prime}}$. Pairs of sites $\{\alpha,\beta\}$ are connected by the space and orbital dependent hopping $T_{ll^{\prime}}(\mathbf{r}_{\alpha\beta})$, see Eq. (A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor).
A striking example of continuously tunable electronic properties are moiré heterostructures—stacks of two-dimensional materials rotated relative to each other by a twist angle [22, 23, 24, 25]. Tuning this angle alters the real-space geometry, which in turn modifies the two components of the wavefunction’s quantum geometric tensor, the quantum metric and Berry curvature [26, 27, 28, 29, 30, 31, 32]. The external control over this tensor has catalyzed the discovery of a wide range of phenomena [33, 25], including superconductivity [23, 34, 35, 36], anomalous and fractional Hall effects [37, 38, 39, 40, 41, 42, 43, 44, 45], Wigner crystallization [46, 47, 48, 49], quasicrystalline behavior [50, 51, 52, 53, 54], and correlated states [22, 55, 56, 57, 58, 59, 60, 61, 62]. However, these remarkable phenomena often occur at specific twist angles, and additionally suffer from twist-angle disorder [63, 64, 65]. Moreover, synthetic platforms like photonic crystals or ultracold atoms cannot be easily twisted. The successes—but also the limitations—of moiré heterostructures therefore motivate the search for new ways to optimize the quantum geometric tensor in solid-state and synthetic matter via real-space geometry.
Here we show that a newly discovered family of aperiodic tilings [66, 67] offers a real-space geometric pathway to tune the quantum geometric tensor. This is the only family of aperiodic tilings parametrically connected by a single real parameter $\ell$  [^1] Aperiodic tilings can be generated by the cut and project method. By varying continuously the cut angle one might expect to realize a continuous family of aperiodic tilings. However, varying this angle continuously typically results in jumps when new tiles appear, see e.g. this video. We thank F. Flicker for this remark.. This parameter determines the position of all atomic sites while preserving the quasicrystalline long-range order. Allowing the hoppings to depend on atomic positions, we show that changes in $\ell$ translate into changes of the quantum geometric tensor [69, 29, 32], expressed in terms of a real-space Chern marker [70] and quantum metric [71, 72]. Changes in the real-space Chern marker induced by $\ell$ induce phase transitions in and out of a topological phase, which exists in an expanded parameter space compared to the same model in crystals. The parameter $\ell$ also tunes the trace of the quantum metric [26, 27, 28, 29, 30], which contributes to the superfluid stiffness and kinetic inductance [73, 74, 75, 25, 76, 77], enters optical sum rules [78, 79, 80], and determines localization [81] and entanglement properties [82]. Hence, changing the real-space geometry using $\ell$ allows one to tune the quantum geometric tensor.
Figure 2: Spectral and topological properties of the model for different parameter values $\ell$. (a) Energy spectrum of the bulk Hamiltonian as a function of the parameter $\ell$. We obtain the bulk spectrum by excluding states whose weight inside the yellow contour in (d) and (e) is smaller than their weight outside it. Magenta dashed lines indicate the $\ell$ values used in panels (b) and (c). The yellow dashed line indicates the Fermi energy $E_{F}$. (b) Spectral function $\mathcal{A}(E,{\bf p})$ for $\ell=0.23$, showing a full gap at the energy indicated by the white dashed line. (c) Spectral function $\mathcal{A}(E,{\bf p})$ for $\ell=1/3$, showing the presence of edge states within bulk gaps. (d) Local Chern marker $\mathcal{C}_{{\bf r}_{i}}$ Eq. (2), evaluated at the energy indicated by the white dashed line in (c) for $\ell=1/3$. The yellow solid line outlines the bulk region used to perform the averaging. In this region, the average value of the local Chern marker is $0.98$. (e) Local quantum metric $g_{\bf r}$, Eq. (5), computed at the $E_{F}$ indicated by the dashed line in (c) for $\ell=1/3$. We observe that the local quantum metric takes large values at the edges, indicating delocalized edge states. In all the panels the system has $2530$ sites and $M/t=-2.7$. In panels (b) and (c) we used the kernel polynomial method [83] to compute the spectral function with 512 moments.
Geometric control of topology—Fig. 1a shows the Hat family of aperiodic tilings, parameterized by a real number $\ell\in\left[0,1\right]$. It parametrizes the two independent lengths of parallel sides, $a=\ell$ and $b=1-\ell$. By tuning $\ell$, one can continuously interpolate between different tiles in the family. The Hat tiling [$\ell=1/(1+\sqrt{3})$] made headlines as the first aperiodic monotile [66, 67]. Other named tiles are the Specter [$\ell=1/2$], the Turtle [$\ell=\sqrt{3}/(1+\sqrt{3})$], the Chevron [$\ell=0$], and the Comet [$\ell=1$] which sets our unit of length. Previous work has shown that a non-bipartite version of the Hat tiling shares similar electron properties with those of graphene [84] when all the hoppings are set to be equal. Here, we focus on the bipartite version of these tilings, and describe their construction in Appendix A of the Supplemental Material (SM) [85], by adapting the code published along with Ref. [86].
Figure 3: Topological phase diagram. (a) Bulk half-signature of the spectral localizer at $E_{0}=E_{F}$ as a function of both the geometrical parameter $\ell$ and the value of the onsite term $M/t$. $E_{F}$ is fixed at $3/4$ filling out of a total of $N_{\mathrm{sites}}\times N_{\mathrm{orbs}}=2530\times 2$ states. We observe topological phases for a wider parameter region compared to the square-lattice crystalline model, whose phase boundaries are delimited by the black dashed lines. (b) Localizer gap corresponding to (a). The gap in the spectrum of the spectral localizer closes at the boundary between the topological and trivial phases of panel (a). (c) Bulk half-signature of the spectral localizer averaged over 50 disorder realizations at $E_{0}=E_{F}$ and $M/t=0.5$, as a function of $\ell$ and onsite disorder strength $W/t$. We observe two types of behavior as we increase $W/t$ for fixed $\ell$: either the clean system is topological and transitions into a trivial Anderson insulator (e.g. at $\ell=0.5$), or a trivial clean system transitions into a topological Anderson insulator, to later become a trivial Anderson insulator (e.g. at $\ell=0.8$ or $\ell=0.2$). (d) The localizer gap vanishes at the interface between the topological and trivial Anderson insulators in (c). In all panels, the system has $N_{\mathrm{sites}}=2530$ sites and $\kappa=0.01$.
To exemplify how $\ell$ can control the electronic quantum geometry we consider a non-crystalline generalization of the Qi-Wu-Zhang Chern insulator model [87] on the vertices of the tiling, see Fig. 1a-e, that reads [88, 89, 90, 91]:
$\mathcal{H}=$ $\sum_{\left\langle\alpha,\beta\right\rangle}\sum_{l,l^{\prime}\in\left\{1,2\right\}}T_{ll^{\prime}}^{\phantom{}}(r_{\alpha\beta}^{\phantom{}},\theta_{\alpha\beta}^{\phantom{}})c^{\dagger}_{\alpha,l}c_{\beta,l^{\prime}}^{\phantom{}}$
$+\sum_{\alpha}\sum_{l,l^{\prime}\in\left\{1,2\right\}}\epsilon_{ll^{\prime}}^{\phantom{}}c_{\alpha,l}^{\dagger}c_{\alpha,l^{\prime}}^{\phantom{}}.$ (1)
The hopping amplitude between two orbitals $l$ and $l^{\prime}$ at sites $\alpha$ and $\beta$ sharing an edge, $T_{ll^{\prime}}(r_{\alpha\beta},\theta_{\alpha\beta})$, depends on the inter-atomic separation $r_{\alpha\beta}=|{\bf r_{\alpha}}-{\bf r_{\beta}}|$ and the polar angle $\theta_{\alpha\beta}$. We choose the hopping to be $T(r_{\alpha\beta},\theta_{\alpha\beta})=h(\theta_{\alpha\beta})f(r_{\alpha\beta})$, with $f(r_{\alpha\beta})=\exp(1-r_{\alpha\beta}/\bar{a})$ and $h(\theta_{\alpha\beta})=-(1/2)\sigma_{z}-(it/2)\left[\cos(\theta_{\alpha\beta})\sigma_{x}+\sin(\theta_{\alpha\beta})\sigma_{y}\right]$, and the onsite energy as $\epsilon=(2+M)\sigma_{z}$. Here $\bar{a}$ denotes the average bond-length and $(\sigma_{x},\sigma_{y},\sigma_{z})$ are the Pauli matrices acting in orbital space. If we would choose $\alpha,\beta$ to lie on a square lattice we would recover the crystalline Chern insulator of Ref. [87].
The corresponding bulk energy spectrum is shown in Fig. 2a as a function of $\ell$. It is symmetric with respect to zero energy as a consequence of a chiral symmetry [90] that places the model is in class $D$ of the Altland-Zirnbauer classification [92]. This feature is not essential to our findings. At the extreme values $\ell\approx 0,1$ the spectrum is gapped and we expect the bands to be topologically trivial. The reason is that at these values of $\ell$ many sites come together in real space, forming clusters of sites (see Fig. 1a and Supplemental Video). The exponential form of $f(r_{\alpha\beta})$ leads to a strong suppression of the hoppings between clusters, exponentially localizing the wave-functions to these clusters. Since exponentially localized wave-functions are topologically trivial, so the gaps are all trivial [93, 94, 95].
This expectation is confirmed by computing the bulk average of the local-Chern marker [96]
$C_{\textbf{r}_{\alpha}}=2\pi\text{Im}\sum_{l}\bra{\textbf{r}_{\alpha},l}\hat{P}\left[\hat{Q}\hat{X},\hat{P}\hat{Y}\right]\hat{P}\ket{\textbf{r}_{\alpha},l},$ (2)
where $\hat{P}=\sum_{E<E_{F}}\ket{E}\bra{E}$ | is the projector onto occupied eigenstates states |$\ket{E}$ below the Fermi energy $E_{F}$, $\hat{Q}=1-\hat{P}$, and $\hat{X}$ and $\hat{Y}$ the position operators in the plane. Its average over a bulk real-space region of area $\mathcal{A}_{b}$ (see Appendix C in SM [85]), $C=(1/\mathcal{A}_{b})\sum_{\alpha\in\text{Bulk}}C_{\textbf{r}_{\alpha}}$, determines the bulk Chern number, $C$, which is approximately quantized if the occupied states are topological [70]. We find that the average of the local-Chern marker is zero when we place $E_{F}$ within any of the gaps seen at $\ell\approx 0,1$, as advertised.
As we tune $\ell$, the angle- and position-dependence of the hoppings in Eq. (A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor) allows the spectrum to evolve, see Fig. 2a. At $\ell\approx 0.28$ we observe band closings, suggesting a topological phase transition.
Topological transitions would imply the appearance of in-gap spectral weight corresponding to topological edge states in the total momentum-resolved density of states, or spectral function. We define the spectral function by projecting the real-space Hamiltonian $\mathcal{H}$ onto plane-wave states, $A(E,\mathbf{p})=\sum_{l}\bra{\textbf{p},l}\delta(\mathcal{H}-E)\ket{\textbf{p},l}$, where |$\ket{\textbf{p},l}=\sum_{\alpha}e^{i\mathbf{p}\mathbf{r}_{\alpha}}\ket{\textbf{r}_{\alpha},l}$, see Appendix B in SM [85]. This function is measurable in angle-resolved photemission spectroscopy even without translational invariance [97, 98, 99, 100, 101, 102]. Note that, unlike the crystalline case, the components of ${\bf p}$ are not restricted to the interval $\left[0,2\pi\right]$. Comparing Figs. 2b and c, at either side of the gap closing (with $\ell=0.23$ and $\ell=0.33$, respectively) we see the appearance of in-gap spectral weight, as expected for a topological transition (see also Supplemental Video).
We confirm the topological nature of the in-gap states at $\ell=0.33$ by computing the local Chern marker at $E_{F}=-1.15$, see Fig. 2d. For these parameters its bulk average is $C\approx 0.98$, confirming the topological phase. These results show that we can tune the topological phase by varying $\ell$ while keeping the filling fixed. For example, at $E_{F}=-1.15$ (dashed line in Fig. 2c), a trivial aperiodic tiling leads to an aperiodic tiling with finite Chern number.
We now ask, how far do the topological properties survive in parameter space, compared to the crystalline lattice model with the same Hamiltonian parameters? Because the bulk average of the local Chern marker is only approximately quantized, it is more convenient to answer this question with a local marker that is exactly an integer in the bulk. Such a marker can be obtained using the spectral localizer [103, 104, 105, 106]
=(^$\mathcal{L}=(\hat{\mathcal{H}}-E_{0}\mathbb{1})\sigma_{z}+\kappa\left[(\hat{X}-x_{0}\mathbb{1})\sigma_{x}+(\hat{Y}-y_{0}\mathbb{1})\sigma_{y}\right],$ (3)
where $E_{0}$ is the reference energy at which we want to probe topology and $x_{0}$ and $y_{0}$ are chosen in the bulk of the system  [^2] In practice, we choose the coordinates $x_{0}$ and $y_{0}$ by computing the geometric mean of the vertex coordinates.. The scalar $\kappa$ is fixed to compensate the differences in scales between $\hat{\mathcal{H}}$ and the position operators $\hat{X}$, and $\hat{Y}$, see Appendix D in SM [85]. The half-signature of the localizer, $\frac{1}{2}\text{Sig}(\mathcal{L})$, defined as the half-difference between the number of positive and negative eigenvalues, is an integer that equals the bulk Chern number [103, 104, 105, 106]. In turn, the minimal eigenvalue of the spectral localizer, the localizer gap, determines the stability of a given phase, since a topological phase transition can only occur if the localizer gap closes [106], see Appendix D in SM [85].
As a reference, when applied on a square lattice, the model defined in Eq. (A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor) only has topological phases when $40<M/t<-$ 4. In contrast, we see in Fig. 3a that the phase diagram of the Hat family as a function of $\ell$ and $M/t$ extends further. The topological bands survive for much larger values of $M/t$ compared to the square lattice model, delimited by the dashed lines in Fig. 3a, especially around $\ell\approx 0.5$. Consistently, we see that the localizer gap, shown in Fig. 3b, closes at the phase boundaries.
Given that $\ell$ serves as a tuning parameter for inducing topological phase transitions, one might ask whether it can also be used to control a disordered induced Anderson transition. To address this question, we introduce a random onsite potential
$V=\sum_{\alpha}\sum_{ll^{\prime}}V_{\alpha,ll^{\prime}}^{\phantom{}}c_{\alpha,l}^{\dagger}c_{\alpha,l^{\prime}}^{\phantom{}},$ (4)
where $V_{\alpha,ll^{\prime}}^{\phantom{}}=V_{\alpha}\delta_{ll^{\prime}}$ represent a site-dependent potential drawn from a uniform distribution in the range $[-W/2,W/2]$, where $W$ controls the disorder strength. We observe in Fig. 3b that the system remains topological for values of $W/t\leq 1$ before undergoing an Anderson transition, signaled by the closing of the average localizer gap, shown in Fig. 3d. As in crystals, for specific values of $\ell$, such as $\ell=0.8$, we observe that the system, initially trivial in the clean limit, transitions into a topological phase as $W$ increases, realizing a topological Anderson insulator  [108, 109, 110]. Interestingly, $\ell$ can be used as an additional knob to control a topological Anderson transition at fixed disorder strength, a mechanism specific to this tiling family.
Figure 4: (a) Energy spectrum as a function of $\ell$, showing a zoomed-in version of Fig. 1a including bulk and edge states. $3/4$ filling is shown as a yellow dashed line in (a). (b) Bulk trace of the local quantum metric $G$ and local Chern marker $C$ for this filling as a function of $\ell$. At small $\ell$ the sites cluster and localize the wave-functions for this filling, leading to a small $G$ with $C\approx 0$. At intermediate $\ell\approx 0.5$, $G\geq C\approx 1$, satisfying the trace inequality. For large $\ell\gtrsim 0.7$ this filling sits in the middle of a band of delocalized states with increased $G$. In both panels, the system has $N_{\mathrm{sites}}=2530$ and $M/t=-2.7$.
Geometric control of the quantum metric—The local Chern marker is the imaginary part of a more general tensor called the quantum geometric tensor [26, 27, 28, 29, 30, 31, 32] , which can also be controlled by tuning $\ell$. Its real part is known as the quantum metric [31, 29, 32], and is proportional to the real-space spread of occupied states. By expressing it as a local real-space quantity (see Appendix C in SM [85] and Refs. [71, 72])
$g_{\bf{r}_{\alpha}}=2\pi\mathrm{Re}\sum_{\mu=\left\{x,y\right\}}\sum_{l}\langle{\bf r}_{\alpha},l\lvert\hat{P}\hat{r}_{\mu}\hat{Q}\hat{r}_{\mu}\hat{P}\rvert{\bf r}_{\alpha},l\rangle,$ (5)
where $\hat{r}_{x}=\hat{X}$ and $\hat{r}_{y}=\hat{Y}$, we can use it to observe edge-state signatures, as shown in Fig. 2e for the topological state at $\ell=0.33$ at $E_{F}=-1.15$. We see $g_{\bf{r}_{\alpha}}$ reaches its maximum values at the edges of the system, as expected for delocalized topological edge states.
As with the bulk local Chern marker, the bulk quantum metric, is also tunable as a function of the parameter $\ell$ at a given filling. We define the bulk quantum metric by averaging Eq. (5) over a real-space region of area $\mathcal{A}_{b}$ as $G=(1/\mathcal{A}_{b})\sum_{\alpha\in\text{Bulk}}g_{{\bf r}_{\alpha}}$, see Appendix C in SM [85]. Tuning $G$ is desirable in several contexts. For instance, $G$ is known to be bounded from below by the Chern number, $G\geq C$, an inequality known as the trace inequality [111, 112, 113, 114, 115, 116]. When a band saturates this inequality it is termed vortexable or ideal [115, 116], as it is a favorable starting point to realize fractional Chern insulator states [117, 118, 119]. The quantum metric enters the superfluid stiffness [73, 74, 75, 25, 76, 77] and optical sum rules bounding topological gaps [111, 78, 79, 80]. The quantum metric depends on the real-space position, or embedding, of the orbitals [120, 121, 122, 123, 124, 125, 31, 126], controlled by $\ell$ in our system. In crystals, the embedding-dependence can be exploited to optimize responses such as the photovoltaic efficiency [124].
Fig. 4 shows the tunability of the quantum geometric tensor by plotting the trace of the local Chern marker $C$ and the local quantum metric $G$ as a function of $\ell$. For 3/4 filling (dashed yellow line in Fig. 4a) $G$ decreases as $\ell$ decreases with $C\approx 0$, see Fig. 4b. This is consistent with the fact that for these parameters the bands are narrow, topologically trivial, and states are localized due to the clustering of sites (see Supplemental Video). In contrast, at intermediate values of $\ell\approx 0.5$ the local Chern marker is approximately quantized to $C\approx 1$. The quantum metric decreases but does not saturate the trace inequality $G\geq C$, as seen in other Chern insulator models [127, 128]. At large $\ell\gtrsim 0.7$, $C$ decreases to zero while $G$ increases. This indicates that the states at this filling are more delocalized, consistent with the fact that $E_{F}$ falls within a trivial band with a relatively large spread in energy, see Fig. 4a.
Conclusions—In this work we have used the real-space geometric parameter $\ell$, which continuously connects members of the Hat family of aperiodic tilings, to tune the quantum geometric tensor. The parametrization in terms of $\ell$ is unique to this family, offering a new knob to control transport properties of aperiodic tilings. By allowing the hopping amplitudes to depend on the relative positions of the atoms, we have shown that $\ell$ controls the transition between trivial and topological phases, also in the presence of disorder. The parameter space supporting topological phases is broader than that of the square-lattice counterpart, highlighting the enhanced tunability offered by these tilings.
We have shown how $\ell$ tunes the quantum metric, which underlies key physical properties such as the superfluid stiffness [73, 74, 75, 76, 25, 129, 77] and the optical weight [130, 79, 131, 132]. The ability to tune the quantum metric suggests avenues to engineer properties of unconventional superconductors [133], photonic systems [134, 135], and correlated electron systems [136, 137, 138, 139].
The continuously parametrization in terms of $\ell$ can be exploited in different physical contexts. The systems we consider here could be realized in electronic systems by lithographic patterning of solid-state systems [140], depositing CO molecules on metals [141, 142], or by engineering 2D Shiba states with magnetic ad-atoms on superconductors [143, 144, 145]. Moiré heterostructures have also been used to realize unconventional tessellations [146]. Choosing different $\ell$ provides a new tuning knob to design topology and geometry in photonic metamaterials [147, 148, 6, 149, 150], polaritonic systems [151], electrical circuits [152], microwave networks [153], acoustic [154] and mechanical [155, 156, 157] metamaterials. More broadly, the tunability of this family of aperiodic tilings has already proven useful in enhancing earthquake detection [86] and could benefit other areas such the study of spin models [158], elasticity [159, 160], and fluids [161] on aperiodic tilings.
Acknowledgments — We thank E. Ben Achour, P. D’Ornellas, F. Flicker, L. Gómez Paz, Q. Marsal, S. Pachhal, M. Thumin, D. Varjas, and M. Yoshii for useful discussions and suggestions. We are grateful to P. D’Ornellas and S. Roche for their critical reading of the manuscript. J. S. is supported by the program QuanTEdu-France n° ANR-22-CMAS-0001 France 2030. A. G. G. is supported by the European Research Council (ERC) Consolidator grant under grant agreement No. 101042707 (TOPOMORPH).
Supplementary Material for “A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor”
Appendix A Generating aperiodic tilings from inflation rules
In this section, we describe the inflation procedure introduced in [66] and employed in [162] to generate the aperiodic tilings we use in the main text. As illustrated in Fig. S1, this procedure involves the successive replacement of one set of tiles with another made of a larger amount of tiles. In this work, we adapt the code used in Ref. [86] to generate the tilings presented in Figs. 1 and 2. We begin by examining a single tile, which is a 14-sided polygon whose edges can be grouped into pairs of equal-length, parallel segments, as shown in Fig. S1a. The first step is to combine one tile with its mirror image, forming a two-tile compound. Following [66], this two-tile compound can then be combined with six additional tiles to create a cluster, referred to as H8, or with five other tiles to form the H7 cluster. The inflation process proceeds as follows: each individual tile is replaced by an H8 cluster, while each two-tile compound is replaced by an H7 cluster. This process can be iterated indefinitely to generate arbitrarily large lattices. The number of tiles after the $n$-th iteration is given by the following recursion relation
$\left(H_{7}\right)_{n+1}$ $=5\left(H_{8}\right)_{n}+\left(H_{7}\right)_{n},$ (6)
$\left(H_{8}\right)_{n+1}$ $=6\left(H_{8}\right)_{n}+\left(H_{7}\right)_{n},$ (7)
where $\left(H_{7}\right)_{n}$ (resp. $\left(H_{8}\right)_{n}$) denotes the number of tiles present in the H7 (resp. H8) cluster after the $n$-th iteration of the inflation procedure. In the main text we present results for the third iteration, which has 2530 sites. We have checked that our results are well converged with system size.
Figure S1: Tile and inflation rule. (a) A tile is a fourteen-sided polygon with edge lengths parameterized by a geometric parameter $\ell$. In this example, the value of $\ell$ is taken to be $1/(1+\sqrt{3})$. (b) To construct the tiling, we use clusters of tiles. We begin by defining a two-tile compound consisting of one tile (in white) and its mirror image (in pale blue). We then establish a substitution rule for both the single tile and the two-tile compound using the H7 and H8 clusters.
Appendix B Bulk and surface spectral function.
In this Appendix we describe explicitly the procedure followed to compute the spectral function shown in Fig. 2. We first introduce the real space Green’s function. For a system composed of $N_{\text{sites}}$ atoms and $N_{\text{orbs}}$ internal degrees of freedom, the real-space Green’s function is an $N_{\text{sites}}N_{\text{orbs}}\times N_{\text{sites}}N_{\text{orbs}}$ matrix which reads
$\mathcal{G}(E)=\lim_{\eta\rightarrow 0}\left(\mathcal{H}-(E-i\eta)\mathbf{1}_{N_{\text{sites}}N_{\text{orbs}}}\right)^{-1},$ (8)
where $\mathcal{H}$ is the real-space Hamiltonian and $E$ is the energy at which the Green’s function is computed. In practice, $\eta$ is taken to be finite. The value of $\eta$ is related to the number of moments in the expansion in terms of Tchebychev’s polynomials used in the kernel polynomial method [83]. We now introduce the tight-binding basis $\lvert{\bf{r}_{\alpha}},l\rangle$ that span the Hilbert-space. Each $\lvert{\bf{r}_{\alpha}},l\rangle$ is localized on a site $\alpha$ and carries an internal orbital degree of freedom labeled $l$. These states are represented as $N_{\text{Sites}}\times N_{\text{Orbs}}$ vectors and satisfy the following relation
$\langle{\bf r}_{\alpha},l^{\prime}\lvert{\bf{r}}_{\beta},l\rangle=\delta_{\alpha\beta}\delta_{ll^{\prime}}.$ (9)
Using these, we introduce the normalized plane-wave basis set
$\lvert{\bf{p}},l\rangle=\frac{1}{\sqrt{N_{\text{sites}}}}\sum_{\alpha}\exp\left(i{\bf p}\cdot{\bf r}_{\alpha}\right)\lvert{\bf r}_{\alpha},l\rangle.$ (10)
One can now compute the total momentum-resolved spectral function defined as
$\mathcal{A}\left(E,{\bf p}\right)=-\frac{1}{\pi}\mathrm{Im}\sum_{l}\langle{\bf p},l\rvert\mathcal{G}(E)\lvert{\bf p},l\rangle=\sum_{l}\langle{\bf{p}},l\rvert\delta(\mathcal{H}-E)\lvert{\bf{p}},l\rangle.$ (11)
Appendix C Quantum geometric tensor in real space
The quantum geometric tensor captures the geometry of quantum states in parameter space. In crystalline systems, Bloch states $\lvert u_{n}(\mathbf{k})\rangle$ form an eigenbasis of the Hamiltonian due to translational symmetry. However, states within the same band are not orthogonal: $\langle u_{n}(\textbf{k}_{1})|u_{n}(\textbf{k}_{2})\rangle\neq\delta(\textbf{k}_{1}-\textbf{k}_{2})$. To quantify how non-orthogonal these states are, Provost and Valle [163] introduced a gauge invariant distance for isolated bands
$d^{(n)}(\textbf{k}_{1},\textbf{k}_{2})=\sqrt{1-|\langle u_{n}(\textbf{k}_{1})|u_{n}(\textbf{k}_{2})\rangle|^{2}}.$ (12)
Considering two states of the same isolated band $n$ distant from $\delta\textbf{k}$, the infintesimal interval $ds^{2}$ can be written as $ds^{2}=\,G^{(n)}_{\mu\nu}(\textbf{k})dk^{\nu}dk^{\nu}$, with
$G^{(n)}_{\mu\nu}(\textbf{k})=\langle\partial_{\mu}u_{n}(\textbf{k})|(1-|u_{n}(\textbf{k})\rangle\langle u_{n}(\textbf{k})|)|\partial_{\nu}u_{n}(\textbf{k})\rangle.$ (13)
$G_{\mu\nu}(\textbf{k})$ is the quantum geometric tensor, and $\partial_{\mu}=\partial/\partial k^{\mu}$ with $\mu,\nu=\{x,y\}$ in two dimensions. Integrating the quantum geometric tensor over the Brillouin zone gives us access to the Hilbert-space volume spanned by the $n$-th state. The real and imaginary parts of the quantum geometric tensor are called the quantum metric and Berry curvature, respectively [96, 70, 69, 29, 32].
Despite the absence of translational symmetries of aperiodic tilings, it is possible to define a bulk, real-space version of the quantum geometric tensor as [164, 69, 29, 32]
$G_{\mu\nu}=\frac{2\pi}{\mathcal{A}_{b}}\sum_{\alpha\in\text{Bulk}}\sum_{l}\langle{\bf r}_{\alpha},l\rvert\hat{P}\hat{{\bf r}}_{\mu}\hat{Q}\hat{{\bf r}}_{\nu}\hat{P}\lvert{\bf r}_{\alpha},l\rangle,$ (14)
where $\mathcal{A}_{b}$ is the bulk area, $\hat{P}$ is the projector over occupied states, $\hat{Q}=1-\hat{P}$ is the projector onto unoccupied states and $\hat{r}_{\mu}$ is defined as
$\hat{r}_{\mu}=\begin{cases}\hat{X}\quad\text{if}\quad\mu=x,\\ \hat{Y}\quad\text{if}\quad\mu=y.\end{cases}$ (15)
To discuss the trace inequality [111, 112, 113, 114, 115, 116], as we do in the main text, it is important to normalize $G_{\mu\nu}$ consistently for all $\ell$. We achieve this by normalizing bulk-averaged quantities over a Voronoi tessellation [165, 166] of the system at each $\ell$, see Fig. S2.
Figure S2: Voronoi tessellation of the lattice obtained for $\ell=0.5$. The Voronoi cells corresponding to bulk sites are highlighted in light blue. The bulk area $\mathcal{A}_{b}$ corresponds to the sum of all the areas corresponding to individual blue cells.
Here, the bulk region is defined via Voronoi cells centered on lattice sites within a finite window, and the bulk area $\mathcal{A}_{b}$ is the sum of their individual areas. By using the Voronoi tessellation we can conveniently account for each of the atomic site’s area that contributes to $\mathcal{A}_{b}$.
The quantum metric computed in the main text is then defined as the sum of the diagonal components of the quantum geometric tensor
$G=\mathrm{Re}\left[G_{xx}+G_{yy}\right].$ (16)
From Eq. (14) one can define a local quantum metric as:
$g_{\bf{r}_{\alpha}}=2\pi\mathrm{Re}\sum_{l}\langle{\bf r}_{\alpha},l\lvert\hat{P}\hat{X}\hat{Q}\hat{X}\hat{P}+\hat{P}\hat{Y}\hat{Q}\hat{Y}\hat{P}\rvert{\bf r}_{\alpha},l\rangle.$ (17)
The quantum metric is then defined as the normalized sum over the bulk of the local quantum metric Eq. (17)
$G=\frac{1}{\mathcal{A}_{b}}\sum_{\alpha\in\text{Bulk}}g_{\bf r_{\alpha}}.$ (18)
The quantum geometric tensor encodes important geometric and topological properties of the electronic structure: The quantum metric quantifies the spatial spread of the occupied states: $G\sim\langle\mathbf{r}^{2}\rangle-\langle\mathbf{r}\rangle^{2}$ and is related to the localization properties of Wannier functions [71, 72].
Analogously, the much more commonly studied local Chern marker is defined by the antisymmetric, imaginary part of the quantum geometric tensor [96, 70]. Its bulk average, as defined in the main text below Eq. (2), converges to the Chern number of filled states, even in systems lacking translational symmetry.
Appendix D Spectral localizer gap and choice of $\kappa$
Figure S3: Localizer gap of the spectral localizer Eq. (19) as a function of the reference energy $E_{0}/t$ and the parameter $\kappa$, measured in units of $t/\ell_{c}$. Here $\ell_{c}=1$ is the side of the Comet tile, see Fig. 1a, and sets our unit length. We compute in the bulk of the system, at a position $x_{0},y_{0}$ chosen by computing the geometric mean of the site’s coordinates, for $\ell=0.5$ and $M/t=0.5$. The vertical dashed black line corresponds to the value of $\kappa$ used in the main text. For $E_{0}/t=2.65$, the energy set by $E_{F}$ at this $\ell$ in Fig. 3, we are within a bubble with a large localizer gap that is well defined for several orders of magnitude in $\kappa$.
In this section, we justify the choice of the parameter $\kappa$ used to generate Fig. 3 via the spectral localizer [103, 104, 105, 106], whose definition we repeat here for convenience
=(^$\mathcal{L}=(\hat{\mathcal{H}}-E_{0}\mathbb{1})\sigma_{z}+\kappa\left[(\hat{X}-x_{0}\mathbb{1})\sigma_{x}+(\hat{Y}-y_{0}\mathbb{1})\sigma_{y}\right].$ (19)
The parameter $\kappa$ balances the typical energy scales of the Hamiltonian $\mathcal{H}$ compared to those of the the position operators $\hat{X}$ and $\hat{Y}$. If $\kappa$ is chosen too small, the localizer $\mathcal{L}$ becomes approximately (^$\left(\hat{\mathcal{H}}-E_{0}\mathbb{1}\right)\sigma_{z}$, whose spectrum is symmetric around zero. Because of this symmetry the signature vanishes, yielding a trivial topological invariant. Conversely, if $\kappa$ is chosen too large, the localizer is dominated by the position operators and effectively decouples from the Hamiltonian, again resulting in a trivial invariant. Therefore, $\kappa$ must be chosen so that the Hamiltonian and the position operators contribute comparably to the localizer, ensuring the computation captures the correct topological features.
To find a suitable value of $\kappa$, it is convenient to compute the localizer gap at different energies for different values of $\kappa$ [106]. The localizer gap is defined as the magnitude of the smallest eigenvalue in absolute value, $\min(\lvert\lambda\rvert)$ where $\lambda\in\mathrm{Sp}(\mathcal{L})$. For a given matrix $A$, $\mathrm{Sp}(A)$ denotes the set of all the eigenvalues of $A$ or spectrum of $A$. Because the invariant can only change when the localizer gap closes, the magnitude of the localizer gap attests to the robustness of a given phase computed for a given choice of parameters $(\kappa,x_{0},y_{0},E_{0})$. A large gap, extended in parameter space, signifies a stable phase [105, 167, 168, 106]. In this sense the localizer gap acts analogously to the band-gap of crystalline systems. In practice, for a given system size there is not a single value of $\kappa$ but rather a window of values of $\kappa$ that capture well the topologically robust properties of the system [106].
In Fig. S3 we plot the localizer gap as a function of both $\kappa$ and $E_{0}/t$ for $x_{0}$ and $y_{0}$ chosen in the bulk of the system, by computing the geometric mean of the site’s coordinates. We observe the emergence of distinct bubbles where the localizer gap is large, similar to other localizer studies [105, 167, 168, 106, 169]. These bubbles correspond to the regions in parameter space where both the Hamiltonian and the position operators enter meaningfully the calculation of the spectral localizer. We have checked that, by selecting a value of $\kappa$ such that the energy of interest lies within one of these bubbles, we ensure that the invariant computed via the spectral localiser at a given energy approximately coincides with the bulk average of the local Chern marker at the same energy. Furthermore, we observe that the bubble at $E_{0}/t=2.65$, the energy set by $E_{F}$ at this $\ell$ in Fig. 3, is particularly extended, which ensures the robustness of the calculation with respect to different choices of the parameter $\kappa$.